# Approximation in Hölder spaces

### v3

@@ 3.1. Lipschitz and smooth approximations
In a real normed space $X$, we consider real-valued functions ``$f"$ of the class $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R})$, the last abbreviated by $\dot{\operatorname{VC}}^{0,\omega}(X).$ By Theorem 1.2 these functions ``$f"$ can be approximated by functions, say ``$g",$ of class $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ with bounded support, in the $\dot{C}^{0,\omega}(X)$ seminorm. Our goal is to further approximate these functions ``$g"$ by Lipschitz functions (whose regularity will vary depending on the smoothness properties of $X$) with bounded support. In approximating ``$g"$ a crucial step is to reduce the problem to the problem of approximating Lipschitz functions by smooth Lipschitz functions with good control on the Lipschitz constants. The main results of this subsection are Theorem 3.3, valid for arbitrary normed spaces, and Theorem 3.5, valid for separable normed spaces $X$ with a fixed degree of smoothness. The proofs of these two Theorems 3.3 and 3.5 employ several technical lemma, which will be reused in the later Subsections 3.2 and 3.3.
@@
The same argument swapping $x$ and $z,$gives $|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\varepsilon\omega(\|x-z\|),$ for every $n\in\mathbb{N}.$ We conclude that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon,$ for $n\geq N.$
For the assertion concerning the boundedness of the supports, we may assume that $f(y)=0,$ whenever $\|y\|\geq R,$ for some $R>0.$ We will localize the infimum defining $f_{n}.$ For each $y\in X,$ it follows from the definition of $f_{n}$, and the fact that $f\in\dot{C}^{0,\omega_{f}}(X),$ that
$f(y)+n\|x-y\|$ $\geq f(x)-\|f\|_{\dot{C}^{0,\omega_{f}}(X)}\omega_{f}(\|x-y\|)+n\|x-y\|$
$\qquad\geq f_{n}(x)-\|f\|_{\dot{C}^{0,\omega_{f}}(X)}\omega_{f}(\|x-y\|)+n\|x-y\|.$ (31)
@@
The inclusion ``"``"`$\subset$"` ` ⊂ " follows immediately from Lemma 3.1. For the converse inclusion ``","```$\supset$",` ` ⊃ " , we observe that condition (3) gives $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$, and then recall that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ is closed.
By Theorem 1.2 we know that if $f\in\dot{\operatorname{VC}}^{0,\omega}(X),$ and $\varepsilon>0,$ we can find $g\in\dot{\operatorname{VC}}^{0,\omega}(X)$ with bounded support and $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Then, Lemma 3.1 provides us with $h\in\operatorname{Lip}_{\operatorname{bs}}(X)$ so that $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and therefore $\|f-h\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon;$ this shows the inclusion ``"."```$\supset$".` ` ⊃ " .
For the converse implication ``"``"`$\subset$"` ` ⊂ ", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X);$ see Remark 2.1. ∎
@@
Let $g$ be as in the assumption, and let $R>0$ be so that $g=0$ outside the ball $B(0,R).$ We first construct a suitable bump that decays from one to zero over the annular region $B(0,\lambda R)\setminus B(0,2R),$ for a certain $\lambda>2$ to be fixed later. Since $X$ admits a $C^{k}\cap\operatorname{Lip}$ bump, by [DevilleGodefroyZizler, Proposition II, 5.1] there exist a constant $10<a<$ 1 and a function $\psi:X\to[0,\infty)$ so that $\psi\in\operatorname{Lip}(X)\cap C^{k}(X\setminus\{0\})$ and
@@
The second term in the sum converges to 00 uniformly on $x.$ For the first term, note that if $\|x\|\leq 2R,$ then $\varphi(x)=1,$ and so $|\varphi(x)-1||g_{n}(x)|=0.$ When $\|x\|\geq 2R,$ we have $\lim_{n}|g_{n}(x)|=|g(x)|=0,$ uniformly on those $x,$ by the bounded support of $g.$ Since also $\varphi$ takes values in $[0,1],$ we conclude that the first term tends uniformly to 00 as $n\to\infty.$ Now we have shown that
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\varepsilon>0.$ By Lemma 3.1 we find an approximation $g\in\operatorname{Lip}(X)$ so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, since $g:X\to\mathbb{R}$ is a Lipschitz function, and $X$ has a $C^{k}\cap\operatorname{Lip}$ bump, as a consequence of a result of Hájek and Johanis [HajJoh10, Corollary 15] (here we use that $X$ is separable), we find a sequence $(g_{n})_{n}$ of $C^{k}\cap\operatorname{Lip}$ functions converging uniformly to $g$ such that $\operatorname{Lip}(g_{n})\leq\lambda\operatorname{Lip}(g),$ for all $n,$ and for some absolute constant $\lambda>0$ that may depend on the space $X.$ Since obviously the function $g$ and the sequence $(g_{n})_{n}$ satisfy all the assumptions of Lemma 3.2, and hence $\lim_{n}\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0,$ completing the proof of the inclusion ``"."```$\supset$".` ` ⊃ " . For the reverse inclusion recall that the assumption $t/\omega(t)\to 0,$ as $t\to 0,$ guarantees that $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X);$ and $\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X)$ is a closed subspace of ${\dot{C}^{0,\omega}(X)}.$
@@
Thus we have shown the inclusion ``"."```$\supset$".` ` ⊃ " . For the converse inclusion ``"``"`$\subset$"` ` ⊂ ", it is enough to observe again (as we did at the end of the proof of Theorem 3.3) that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of ${\dot{C}^{0,\omega}(X)},$ again, by Remark 2.1.
@@
[…]
@@
    $k(p)=\begin{cases}\infty\qquad&\text{ if }p\text{ is even},\\ p-1,\qquad&\text{ if }p\text{ is odd},\\ \left\lfloor p\right\rfloor,\qquad&\text{ if }p\text{ is not an integer}.\end{cases}$
@@ 3.2. Approximation in super-reflexive spaces
[…]
@@
[…]
@@
for all $y,z\in E,$ $x\in X;$ indeed, $G(y)=G(z)=0$ is the zero functional, and for $f(y)-f(z),$ by symmetry, it is enough to verify the case $y\in S$ and $\operatorname{\mathit{d}}(z,S)\geq 1$ and $x\in X$ arbitrary, but this case is clear by inspection, since
@@
By the comment just after formula (35), the function $u\mapsto\psi(u):=-\|u\|^{1+\alpha}$ is $C^{1,\alpha}$, and their Fréchet derivative $D\psi$ satisfies $\|D\psi\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq B(\alpha,C).$ Thus, using the Fundamental Theorem of Calculus and some elementary computations, $\psi$ is $At^{1+\alpha}$-strongly paraconvex for some $A=A(\alpha,C);$ see the argument near the end of [AzaMud21, the proof of Lemma 3.6]. It turns out that then both $m$ and $(-g)$ are $AM^{*}t^{\alpha}$-strongly paraconvex. Also, $m\leq g$ on $X,$ and these properties permit to prove the regularity $F\in C^{1,\alpha}(X)$, along with (36) and that $(F,DF)=(f,G)$ on $E.$ See [AzaMud21, Theorems 1.9 and 4.1] for further explanations and details of these concepts and their proofs. Now, by the definition of $F,$ the pointwise relation
@@
holds true. Since $G\equiv 0$ and $\|f\|_{\infty}=1,$ the definition of $m$ and $g$ and this estimate imply that $-2\leq F\leq 2$ on $X.$ It only remains to slightly modify $F$ so that the final function $h_{S}$ takes values in $[0,1].$
@@
[…]
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\varepsilon>0.$ Applying Lemma 3.1, we can find a function $g\in\operatorname{Lip}(X)$ so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, by Theorem 3.8(2), there is a sequence $(g_{n})$ of $\operatorname{Lip}(X)\cap C^{1,\alpha}(X)$ functions with $\sup_{n}\operatorname{Lip}(g_{n})\leq\kappa\operatorname{Lip}(g),$ and $\lim_{n}\sup_{X}|g-g_{n}|=0.$ Therefore Lemma 3.2 yields that $\lim_{n}\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0,$ and taking some $g_{n}$ from the sequence with $\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ we conclude the inclusion ``""```$\supset$"` ` ⊃ " of $(i).$ The reverse inclusion follows by noting that $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$ by the condition (3), and that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X).$
@@ 3.3. Approximation of Banach-valued mappings from $c_{0}$
where $\lambda_{|F|}$ is the Lebesgue measure on $\mathbb{R}^{|F|}.$
@@
Since $g\in\dot{C}^{0,\omega}(X,Y)$ and $P_{S_{x}}$ is $1$-Lipschitz, the function $z\mapsto(g\circ P_{S_{x}})(z)$ belongs to $\dot{C}^{0,\omega}(X,Y)$, implying that $z\mapsto(g\circ P_{S_{x}})(z)$ is uniformly continuous on $X.$ Therefore, by the previous formula, $B(x,r)\ni z\mapsto h_{S_{x}}(z)$ is a finite dimensional smooth mollification (in the Bochner sense) with a uniformly continuous mapping, which shows that $h_{S_{x}}\in C^{\infty}(B(x,r),Y).$ Moreover, using Fubini’s theorem, one can easily check that if $F\subset\mathcal{A}$ is finite and $F\supset S_{x},$ then also $h_{F}=h_{S_{x}}$ on $B(x,r).$ This enables us to define the desired mapping $h:X\to Y$ in the following manner. Given $x\in X,$ let $S_{x}\subset\mathcal{A}$ be the subset described above, and define $h(x):=h_{S_{x}}(x).$ Moreover, ordering the collection of finite subsets $\mathcal{F}(\mathcal{A})$ of $\mathcal{A}$ by inclusion, the pointwise limit of the net $\{h_{S}\}_{S\in\mathcal{F}(\mathcal{A})}$ is exactly $h.$
@@
For the particular case of $\mathcal{A}=\{1,\ldots,n\},$ $c_{0}(\mathcal{A})$ becomes $\mathbb{R}^{n}$ with the supremum norm, which is of course equivalent with the usual Euclidean setting. Theorem 3.12 then has the following consequence.

### v4

@@ 3.1. Lipschitz and smooth approximations
In a real normed space $X$, we consider real-valued functions f``f" of the class $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R})$, the last abbreviated by $\dot{\operatorname{VC}}^{0,\omega}(X).$ By Theorem 1.2 these functions f``f" can be approximated by functions, say g,``g", of class $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ with bounded support, in the $\dot{C}^{0,\omega}(X)$ seminorm. Our goal is to further approximate these functions g``g" by Lipschitz functions (whose regularity will vary depending on the smoothness properties of $X$) with bounded support. In approximating g``g" a crucial step is to reduce the problem to the problem of approximating Lipschitz functions by smooth Lipschitz functions with good control on the Lipschitz constants. The main results of this subsection are Theorem 3.3, valid for arbitrary normed spaces, and Theorem 3.5, valid for separable normed spaces $X$ with a fixed degree of smoothness. The proofs of these two Theorems 3.3 and 3.5 employ several technical lemma, which will be reused in the later Subsections 3.2 and 3.3.
@@
The same argument swapping $x$ and z,z,gives $|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\varepsilon\omega(\|x-z\|),$ for every $n\in\mathbb{N}.$ We conclude that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon,$ for $n\geq N.$
For the assertion concerning the boundedness of the supports, we may assume that $f(y)=0,$ whenever $\|y\|\geq R,$ for some $R>0.$ We will localize the infimum defining $f_{n}.$ For each $y\in X,$ it follows from the definitions of $f_{n}$ and $\omega_{f}$ that
$f(y)+n\|x-y\|$ $\geq f(x)-\omega_{f}(\|x-y\|)+n\|x-y\|$
$\geq f_{n}(x)-\omega_{f}(\|x-y\|)+n\|x-y\|.$ (31)
Using property (30) of the minimal modulus of continuity $\omega_{f}$, and the observations subsequent to the definition of $\omega_{f},$ we have that
$\omega_{f}(\|x-y\|)\leq 2\|x-y\|\omega_{f}(1)\leq 2\omega(1)\|f\|_{\dot{C}^{0,\omega}(X)}\|x-y\|,\quad\text{when}\quad\|x-y\|\geq 1.$
@@
The inclusion $\subset$" follows immediately from Lemma 3.1. For the converse inclusion $\supset$", we observe that condition (3) gives $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$, and then recall that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ is closed.
By Theorem 1.2 we know that if $f\in\dot{\operatorname{VC}}^{0,\omega}(X),$ and $\varepsilon>0,$ we can find $g\in\dot{\operatorname{VC}}^{0,\omega}(X)$ with bounded support and $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Then, Lemma 3.1 provides us with $h\in\operatorname{Lip}_{\operatorname{bs}}(X)$ so that $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and therefore $\|f-h\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon;$ this shows the inclusion $\supset$".
For the converse implication $\subset$", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X);$ see Remark 2.1. ∎
@@
Let $g$ be as in the assumption, and let $R>0$ be so that $g=0$ outside the ball $B(0,R).$ We first construct a suitable bump that decays from one to zero over the annular region $B(0,\lambda R)\setminus B(0,2R),$ for a certain $\lambda>2$ to be fixed later. Since $X$ admits a $C^{k}\cap\operatorname{Lip}$ bump, by [DevilleGodefroyZizler, Proposition II, 5.1] there exist a constant $0<a<1$ and a function $\psi:X\to[0,\infty)$ so that $\psi\in\operatorname{Lip}(X)\cap C^{k}(X\setminus\{0\})$ and
@@
The second term in the sum converges to 0 uniformly on $x.$ For the first term, note that if $\|x\|\leq 2R,$ then $\varphi(x)=1,$ and so $|\varphi(x)-1||g_{n}(x)|=0.$ When $\|x\|\geq 2R,$ we have $\lim_{n}|g_{n}(x)|=|g(x)|=0,$ uniformly on those $x,$ by the bounded support of $g.$ Since also $\varphi$ takes values in $[0,1],$ we conclude that the first term tends uniformly to 0 as $n\to\infty.$ Now we have shown that
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\varepsilon>0.$ By Lemma 3.1 we find an approximation $g\in\operatorname{Lip}(X)$ so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, since $g:X\to\mathbb{R}$ is a Lipschitz function, and $X$ has a $C^{k}\cap\operatorname{Lip}$ bump, as a consequence of a result of Hájek and Johanis [HajJoh10, Corollary 15] (here we use that $X$ is separable), we find a sequence $(g_{n})_{n}$ of $C^{k}\cap\operatorname{Lip}$ functions converging uniformly to $g$ such that $\operatorname{Lip}(g_{n})\leq\lambda\operatorname{Lip}(g),$ for all $n,$ and for some absolute constant $\lambda>0$ that may depend on the space $X.$ Since obviously the function $g$ and the sequence $(g_{n})_{n}$ satisfy all the assumptions of Lemma 3.2, and hence $\lim_{n}\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0,$ completing the proof of the inclusion $\supset$". For the reverse inclusion recall that the assumption $t/\omega(t)\to 0,$ as $t\to 0,$ guarantees that $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X);$ and $\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X)$}.
@@
Thus we have shown the inclusion $\supset$". For the converse inclusion $\subset$", it is enough to observe again (as we did at the end of the proof of Theorem 3.3) that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X)$}, again, by Remark 2.1.
@@
    If $X$ is a separable Banach space with an equivalent norm of class $C^{k}$, then Theorem 3.5 applies for $X$ and $k.$ Indeed, denote such an equivalent norm by $\psi:X\to\mathbb{R}$; and we can assume that $\psi$ satisfies the inequalities on the line (32). As $\psi\in C^{k}(X\setminus\{0\}),$ and since $\psi$ is subadditive in $X,$ it is clear that $\psi$ is $1$-Lipschitz (with respect to the original norm $\|\cdot\|$) on $X.$ If we now pick $\theta:\mathbb{R}\to[0,1]$ of class $C^{\infty}\cap\operatorname{Lip}$, with $\theta=1$ on $(-\infty,1/2]$ and $\theta=0$ on $1,+\infty),$ it is easy to see that the composition $\theta\circ\psi$ defines a $C^{k}\cap\operatorname{Lip}$ bump function on $X.$ Thus, we are in assumptions of Theorem 3.5 for $X$ and $k.$
@@
[…]
@@ 3.2. Approximation in super-reflexive spaces
Let us now approximate real-valued functions of the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}(X)$ over a super-reflexive Banach space $X.$ We remind that $X$ is super-reflexive if every Banach space $Y$ that is finitely representable into $X$ is reflexive. According to Pisier’s renorming theorem [Pisier], a Banach space $X$ is super-reflexive if and only if $X$ admits a renorming (which we keep denoting by $\|\cdot\|$), an exponent $\alpha\in(0,1]$, and a constant $C>0$ for which
@@
Let $X$ be a super-reflexive space, let $\alpha\in(0,1]$ and $C>0$ be so that a renorming of $\|\cdot\|$ of $X$ satisfies (35). Then, there exists a constant $\kappa\geq 1$ depending only on $\alpha$ and $C$ for which the following hold.
@@
for all $y,z\in E,$ $x\in X;$ indeed, $G(y)=G(z)=0$ is the zero functional, and for f(y)f(z),f(y)-f(z), by symmetry, it is enough to verify the case $y\in S$ and $\operatorname{\mathit{d}}(z,S)\geq 1$ and $x\in X$ arbitrary, but this case is clear by inspection, since
@@
By the comment just after formula (35), the function $u\mapsto\psi(u):=-\|u\|^{1+\alpha}$ is $C^{1,\alpha}$, and their Fréchet derivative $D\psi$ satisfies $\|D\psi\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq B(\alpha,C).$ Thus, using the Fundamental Theorem of Calculus and some elementary computations, $\psi$ is $At^{1+\alpha}$-strongly paraconvex for some $A=A(\alpha,C);$ see the argument near the end of [AzaMud21, the proof of Lemma 3.6]. It turns out that then both $m$ and (g)(-g) are $AM^{*}t^{\alpha}$-strongly paraconvex. Also, $m\leq g$ on $X,$ and these properties permit to prove the regularity $F\in C^{1,\alpha}(X)$, along with (36) and that $(F,DF)=(f,G)$ on $E.$ See [AzaMud21, Theorems 1.9 and 4.1] for further explanations and details of these concepts and their proofs. Now, by the definition of $F,$ the pointwise relation
@@
holds true. Since $G\equiv 0$ and $\|f\|_{\infty}=1,$ the definition of $m$ and $g$ and this estimate imply that $2\leq F\leq 2$ on $X.$ It only remains to slightly modify $F$ so that the final function $h_{S}$ takes values in $[0,1].$
@@
Let $X$ be a super-reflexive space, and let $\alpha\in(0,1]$ and $C>0$ be so that a renorming of $\|\cdot\|$ of $X$ satisfies (35). Then the following hold:
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\varepsilon>0.$ Applying Lemma 3.1, we can find a function $g\in\operatorname{Lip}(X)$ so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, by Theorem 3.8(2), there is a sequence $(g_{n})$ of $\operatorname{Lip}(X)\cap C^{1,\alpha}(X)$ functions with $\sup_{n}\operatorname{Lip}(g_{n})\leq\kappa\operatorname{Lip}(g),$ and $\lim_{n}\sup_{X}|g-g_{n}|=0.$ Therefore Lemma 3.2 yields that $\lim_{n}\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0,$ and taking some $g_{n}$ from the sequence with $\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ we conclude the inclusion $\supset$" of $(i).$ The reverse inclusion follows by noting that $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$ by the condition (3), and that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X).$
@@ 3.3. Approximation of Banach-valued mappings from $c_{0}$
where $\lambda_{|F|}$ is the Lebesgue measure on $\mathbb{R}^{|F|}.$ That the integral $h_{F}(x)$ is well-defined for every $x\in X$ can be justified with the same arguments we used in the proof of Theorem 1.6. Namely, if $Q:=[-\eta,\eta]^{|F|},$ the function
$Q\ni t\mapsto\sigma_{x}(t):=g\Big(x-\sum_{\alpha\in F}t_{\alpha}e_{\alpha}\Big)\prod_{\alpha\in F}\theta(t_{\alpha})$
is uniformly continuous $Q\to Y$, by the compactness of $Q$ and the continuity of the functions $g$ and $\theta.$ For each $y^{*}\in Y^{*},$ the mapping $Q\to\mathbb{R}$ given by $Q\ni t\mapsto y^{*}(\sigma_{x}(t))$ is continuous, and thus $\lambda_{|F|}$-measurable over $Q.$ By continuity of $\sigma_{x}$, the image $\sigma_{x}(Q)$ is compact, and so $\overline{\mathrm{span}}(\sigma_{x}(Q))$ is separable. By [BenyaminiLindenstrauss, Proposition 5.1], this shows the measurability of $\sigma_{x}$ in $Q.$ Also, by [BenyaminiLindenstrauss, Proposition 5.2] $\sigma_{x}$ is Bochner integrable if $\int_{Q}\|\sigma_{x}(t)\|_{Y}\,\mathrm{d}\lambda_{|F|}(t)<\infty,$ which is true because $\theta$ is a test function, and $g,$ being in $\dot{C}^{0,\omega}(X,Y),$ is bounded on bounded sets.
@@
Since $g\in\dot{C}^{0,\omega}(X,Y)$ and $P_{S_{x}}$ is $1$-Lipschitz, the function $z\mapsto(g\circ P_{S_{x}})(z)$ belongs to $\dot{C}^{0,\omega}(X,Y)$, implying that $z\mapsto(g\circ P_{S_{x}})(z)$ is uniformly continuous on $X.$ Therefore, by the previous formula, $B(x,r)\ni z\mapsto h_{S_{x}}(z)$ is a finite dimensional smooth mollification (in the Bochner sense) with a uniformly continuous mapping. As we explained in the proof of Theorem 1.6, this guarantees that $h_{S_{x}}\in C^{\infty}(B(x,r),Y).$ Moreover, using Fubini’s theorem, one can easily check that if $F\subset\mathcal{A}$ is finite and $F\supset S_{x},$ then also $h_{F}=h_{S_{x}}$ on $B(x,r).$ This enables us to define the desired mapping $h:X\to Y$ in the following manner. Given $x\in X,$ let $S_{x}\subset\mathcal{A}$ be the subset described above, and define $h(x):=h_{S_{x}}(x).$ Moreover, ordering the collection of finite subsets $\mathcal{F}(\mathcal{A})$ of $\mathcal{A}$ by inclusion, the pointwise limit of the net $\{h_{S}\}_{S\in\mathcal{F}(\mathcal{A})}$ is exactly $h.$
@@
For the particular case of $\mathcal{A}=\{1,\ldots,n\},$ $c_{0}(\mathcal{A})$ becomes $\mathbb{R}^{n}$ with the supremum norm, which is of course equivalent with the usual Euclidean setting. In $\mathbb{R}^{n}$ functions with bounded support have compact support and thus Theorem 3.12 has the following consequence.